1.3 What is f o g if:
f(x) = -2x+7 and g(x)=x+2
Thank you for your assistance, I am rather lost on this problem.

Answer by oscargut(2103)

f o g = f(g(x))
Then
f o g = f(x+2)= -2(x+2)+7 = -2x+3
If you need more help you can contact me at
mthman@gmail.com
Thanks